OpenStudy (anonymous):

Tamara has decided to start saving for spending money for her first year of college. Her money is currently in a large suitcase under her bed, modeled by the function s(x) = 450. She is able to babysit to earn extra money and that function would be a(x) = 6(x − 2), where x is measured in hours. Explain to Tamara how she can create a function that combines the two, and describe any simplification that can be done.

OpenStudy (anonymous):

add the two functions together s(x) + a(x) = 450 + 6(x - 2)

OpenStudy (anonymous):

That's it? How would I explain that?

OpenStudy (anonymous):

that's all I would do

OpenStudy (anonymous):

Thanks! Can u answer anotther?

OpenStudy (anonymous):

Brennan has been playing a game where he can create towns and help his empire expand. Each town he has allows him to create 1.15 times as many villagers. The game gave Brennan 5 villagers to start with. Help Brennan expand his empire by solving for how many villagers he can create with 15 towns. Then explain to Brennan how to create an equation to predict the number of villagers for any number of towns. Show your work and use complete sentences.

OpenStudy (anonymous):

You need to write the formula for a geometric sequence because it's increasing at the constant rate of 1.15. \[a_n=ar^{n-1}\] a = first term = 5 r = common ratio = 1.15

OpenStudy (anonymous):

OK:)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Then do I just plug it in?

OpenStudy (anonymous):

isthe answer 8.74

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

no probleme

OpenStudy (anonymous):

hey i need to try and help those in need